# HyperSHAP: Shapley Values and Interactions for Hyperparameter Importance

### v1

@@ 5 Experiments
In this section, we evaluate the effectiveness and applicability of HyperSHAP across various scenarios. The experiments demonstrate how our games help explain HPIs, interactions, and biases in HPO. In all experiments we rely on four HPO benchmarks; lcbench (Zimmer et al., 2021), rbv2_ranger (Pfisterer et al., 2022), PD1 (Wang et al., 2024), and JAHS-Bench-201 (Bansal et al., 2022). The implementation is based around shapiq (Muschalik et al., 2024a). For more details regarding the experimental setup, we refer to Appendix B. Additional results are contained in Appendix E and a guide for interpreting interaction-based visualizations is presented in Appendix C. Generally, positive interactions are colored in red and negative in blue.
Figure 3: Left: Interaction graphs showing Möbius interactions (MI), Shapley interactions (SI), and Shapley values (SV) where pure MIs are summarized for improved interpretability via SIs and SVs. Right: Faithfulness of the lower-order explanations approximating higher-order effects (Muschalik et al., 2024a). An explanation order of 3 already closely approximates the full game ($R^{2}\approx 1$).
@@ 5.1 Insights from Ablation and Tunability
First, we compare the results of the Ablation and the Tunability game in terms of hyperparameter importance and interactions (cf. Figure 2). We retrieve an optimized configuration of PD1’s lm1b_transformer scenario and explain it with the Ablation game. HyperSHAP’s explanation shows that the majority of the performance increase is attributed to the hyperparameter L-I, which is not surprising since the initial learning rate is also intuitively the most important one. However, using HyperSHAP to create Tunability explanations reveals that both hyperparameters, L-I and O-M (optimizer momentum), are of equal importance with a negative interaction. This means that the optimizer chose to tune L-I over O-M for the configuration in question, even though a similar performance improvement could have been achieved by tuning O-M instead. Hence, HyperSHAP is able to reveal which hyperparameters were subject to optimization via the Ablation game, while the Tunability game emphasizes the potential contributions of hyperparameters together with the interactions between them.
@@ 5.2 Higher-Order Interactions in HPO
Second, we investigate the interaction structure of the HPO problem. In Figure 3, left (MI), and further in Appendix E, we observe the presence of many higher-order interactions, which are difficult to interpret. The SIs (order 2) and SV in HyperSHAP summarize the MI into interpretable explanations. Yet, Figure 3, right, shows that SIs still faithfully capture the overall game behavior, which we measure with a Shapley-weighted loss (Muschalik et al., 2024a) and varying explanation order (cf. Appendix D). We find that most of the explanatory power is captured by interactions up to the third order, confirming prior research that suggests hyperparameter interactions are typically of lower order (Pushak & Hoos, 2020). Interactions beyond the third order contribute little to the overall understanding of the game. Thanks to the convenient properties of the SV and SIs, HyperSHAP provides a reliable way to capture and fairly summarize higher-order interactions into more interpretable explanations.
Figure 5: HyperSHAP explains the surrogate model in SMAC’s Bayesian optimization at 1%, 5%, 25%, and 100% of the budget.
@@ 5.3 Detecting Optimizer Bias
The third experiment uses the Optimizer Bias game to uncover issues in black-box hyperparameter optimizers. To this end, we create two biased hyperparameter optimizers. The first optimizer tunes each hyperparameter separately, ignoring interactions between them, while the second is not allowed to tune the most important hyperparameter. Ideally, a perfect optimizer would show no interactions and no main effects in HyperSHAP’s Optimizer Bias explanations.
@@
Figure 4 shows the Optimizer Bias explanations, i.e., the difference between two Tunability games, using the optimizer’s returned value and the maximum, respectively. Since the main effects are always positive, the differences highlight the optimizer’s inability to properly tune certain hyperparameters. In Figure 4a, small main effects suggest that the optimizer can effectively tune hyperparameters individually, but the presence of both negative and positive interactions shows that it fails to capture these dependencies. This confirms the expectation that this optimizer, which tunes hyperparameters separately, fails to exploit synergies in tuning hyperparameters jointly. On the other hand, the second optimizer, ignoring the weight decay (W-D) hyperparameter for this particular dataset, clearly demonstrates bias in the interaction graph in Figure 4b. The blue main effect for W-D and interactions involving W-D reveal this bias, showing how HyperSHAP can help identify such flaws and contribute to the development of more effective HPO methods.
5.4 Explaining the Surrogate During Optimization
Inspired by Rodemann et al. (2024), we explain SMAC (Lindauer et al., 2022), a state-of-the-art hyperparameter optimizer based on Bayesian optimization, using HyperSHAP to analyze its surrogate model during the optimization procedure. We run SMAC (as pure black-box optimizer) with a budget of around 6 00060006\,0006 000 evaluations and investigate the surrogate model at $1\%,5\%,25\%,$ and $100\%$ of the budget. In Figure 5, we show how the model’s belief about HPI and interactions evolves. Initially uncertain, indicated by large interactions between hyperparameters, SMAC’s surrogate model successfully identifies the most important hyperparameters early on but requires a lot more evaluations to eventually learn the interaction effects correctly. This demonstrates the potential of HyperSHAP to better understand what actually happens in HPO.
@@ 5.4 Explaining the Surrogate During Optimization
Figure 6: Anytime performance plots of hyperparameter optimization runs involving only the top-2 important hyperparameters for two datasets of lcbench (Zimmer et al., 2021).
5.5 Comparison with Functional ANOVA (fANOVA)
Lastly, we design a validation task for hyperparameters deemed important by HyperSHAP, and compare it to the well-known fANOVA framework (Hutter et al., 2014). To this end, for a given HPO task, we run fANOVA and HyperSHAP using the SV with the Tunability and Sensitivity game to calculate the HPI of order 1. We then select the two most important hyperparameters and conduct an HPO run on a reduced HPC space containing only these two hyperparameters. We expect that HyperSHAP achieves better performance for both the Tunability and Sensitivity game since the SV additionally reflects higher-order interactions. Moreover, the hyperparameters chosen based on the Tunability game are expected to outperform the selection based on the Sensitivity game. The results shown in Figure 6 (and further in Section E.2) confirm that the anytime performance of the runs informed by HyperSHAP are clearly superior to those informed by fANOVA, and that Tunability outperforms Sensitivity. This demonstrates that the HPI scores provided by HyperSHAP are effectively contributing to maximizing the accuracy and thus guiding the optimization process. Furthermore, the optimization runs informed by fANOVA converge earlier than those of HyperSHAP, supporting its capability to provide more actionable insights for HPO.
@@ 6 Conclusion
In this paper, we proposed HyperSHAP, a post-hoc explanation framework for consistently and uniformly quantifying HPI using the SV and SIs across three levels: hyperparameter values, tunability of learners, optimizer capabilities. Unlike previous methods that quantify variance (Hutter et al., 2014; Watanabe et al., 2023), HyperSHAP attributes performance contributions. We demonstrated that HyperSHAP not only enhances understanding of the impact of hyperparameter values or tunability of learners but also provides actionable insights for subsequent HPO runs.
Limitations and Future Work.
While HyperSHAP reveals interesting insights into HPO, it is not directly clear how to utilize its explanations for the HPO process itself. This offers room for interesting future work as explanations can be computed in parallel and potentially guide the HPO tuner’s or AutoML framework’s search. Further, the computational bottleneck is the approximation of the $\arg\max$, requiring research on more efficient yet unbiased methods. In future work, we aim to extend the framework to combined algorithm selection and HPO, as well as the design of complete ML pipelines (Olson & Moore, 2016; Wever et al., 2018; Heffetz et al., 2020; Feurer et al., 2022). While these more complex AutoML scenarios are less explored, HyperSHAP provides a versatile and theory-grounded approach for in-depth study. Additionally, we plan to develop HPO methods that utilize HPI, with the goal of learning across datasets to improve optimizer efficiency. This may allow warm-starting HPO in an interpretable way, complementing recent work on prior-guided HPO (Hvarfner et al., 2024) and human-centered AutoML (Lindauer et al., 2024).
Impact Statement
In conducting this research on HyperSHAP, we have carefully considered the ethical implications of our work. This paper presents work with the goal to advance the field of machine learning (ML) and specifically the field of explainable artificial intelligence (XAI) and hyperparameter optimization (HPO). There are many potential societal consequences of our work. The aim of our study is to improve the transparency and interpretability of hyperparameter optimization, which is crucial for building trust and accountability into hyperparameter optimization methods. Thus, our research impacts a wide variety of ML application domains and therein can positively impact ML adoption and potentially reveal biases or unwanted behavior in HPO systems.
However, we recognize that the increased explainability provided by XAI also carries ethical risks. There is the potential for “explainability-based white-washing”, where organizations, firms, or institutions might misuse XAI to justify questionable actions or outcomes. With responsible use, XAI can amplify the positive impacts of ML, ensuring its benefits are realized while minimizing harm.
@@ Appendix A Proofs
$\nu(S)=\max_{\bm{\lambda}\in\bm{\Lambda}}\,\textsc{Val}_{u}(\bm{\lambda}\oplus_{S}\bm{\lambda}^{0},D)=\max_{\bm{\lambda}^{\ast}\in A}\textsc{Val}_{u}(\bm{\lambda}^{\ast},D)\overset{A\subseteq B}{\leq}\max_{\bm{\lambda}^{\ast}\in B}\textsc{Val}_{u}(\bm{\lambda}^{\ast},D)=\max_{\bm{\lambda}\in\bm{\Lambda}}\,\textsc{Val}_{u}(\bm{\lambda}\oplus_{T}\bm{\lambda}^{0},D)=\nu(T).$
This concludes that the Tunability game is monotone. As a consequence, we obtain non-negative SVs due to the monotonicity axiom (Fujimoto et al., 2006) of the SV. We can also give a direct proof of this via the well-known representation of the SV in terms of a weighted average over marginal contributions as
@@
Moreover, the pure individual (main) effects obtained from the functional ANOVA framework are represented by the MI of the individuals (Fumagalli et al., 2024a). By the monotonicity of $\nu$, we obtain again
@@
Given the synthetic Tunability and Sensitivity game with two dimensions $\mathcal{N}=\{1,2\}$, our goal is to show that the main and interaction effects are given by Table 1.
@@
We first proceed to compute the game values of the Tunability game for $S\subseteq\mathcal{N}$ with the optimal HPC $\bm{\lambda}^{\ast}=(1,m)$ as
$\nu_{G_{T}}(S)=\max_{\bm{\lambda}\in\bm{\Lambda}}\textsc{Val}_{u}(\bm{\lambda}\oplus_{S}\bm{\lambda}^{0},D)=\textsc{Val}_{u}(\bm{\lambda}^{*}\oplus_{S}\bm{\lambda}^{0},D)=\sum_{i\in\{1,2\}}\mathbf{1}_{(\bm{\lambda}\oplus_{S}\bm{\lambda}^{0})_{i}=\lambda^{\ast}_{i}}=|S|+\sum_{i\in\mathcal{N}:i\notin S}\mathbf{1}_{\lambda^{0}_{i}=\lambda^{\ast}_{i}}.$ (1)
For the baseline set to $\bm{\lambda}^{0}:=(0,0)$, the second sum in Equation 1 vanishes and we thus obtain
@@
Clearly, the interaction $\lambda_{1}\times\lambda_{2}$, i.e., $m(\{1,2\})$, vanishes, and thus the HPI scores of the individuals are given by their main effects in terms of the MIs. In summary, the HPI main effects using the SV and the MI are both equal to $1$, whereas the interaction is zero, confirming the values shown in Table 1.
For the baseline set to $\bm{\lambda}^{0}:=\bm{\lambda}^{\ast}$, the second sum in Equation 1 equals $|\mathcal{N}|-|S|$ and thus we obtain a constant game
@@
Consequently, all interactions and main effects are zero due to the dummy axiom (Fujimoto et al., 2006), confirming Table 1.
@@
$\nu_{G_{V}}(S)=\mathbb{V}_{\bm{\lambda}^{\ast}\sim p(\bm{\lambda}^{\ast})}[\textsc{Val}_{u}(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0},D)]=\mathbb{V}_{\bm{\lambda}^{\ast}\sim p(\bm{\lambda}^{\ast})}[\sum_{i\in\{1,2\}}\mathbf{1}_{(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0})_{i}=\lambda^{0}_{i}}]=\sum_{i\in S}\mathbb{V}_{\lambda_{i}^{\ast}\sim p(\lambda_{i}^{\ast})}[\mathbf{1}_{(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0})_{i}=\lambda^{0}_{i}}]$ (2)
@@
Given any baseline, we have $\mathbf{1}_{(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0})_{i}=\lambda^{0}_{i}}\sim\text{Ber}(q_{i})$ with $q_{1}=1/2$ and $q_{2}=1/(m+1)$ due to the uniform distribution, which sets this value to $1$, if the optimal HPC value is chosen. The variance of this Bernoulli variable is then given by $q(1-q)$, which yields
@@
$\nu_{G_{V}}(S)=\begin{cases}0,&\text{ if }S=\emptyset,\\ \frac{1}{4},&\text{ if }S=\{1\},\\ \frac{m}{(m+1)^{2}},&\text{ if }S=\{2\},\\ \frac{1}{4}+\frac{m}{(m+1)^{2}},&\text{ if }S=\{1,2\}.\end{cases}$
@@
$m_{G_{V}}(S)=\sum_{L\subseteq S}(-1)^{|S|-|L|}\nu_{G_{V}}(L)=\begin{cases}0,&\text{ if }S=\emptyset,\\ \frac{1}{4},&\text{ if }S=\{1\},\\ \frac{m}{(m+1)^{2}},&\text{ if }S=\{2\},\\ 0,&\text{ if }S=\{1,2\},\end{cases}$
which confirms the values given in Table 1 and concludes the proof. ∎
A.3 Example: Non-Monotone Sensitivity Game

### v2

@@ 6 Experiments
We evaluate the applicability of HyperSHAP across various explanation tasks and benchmarks. To this end, we rely on four HPO benchmarks: lcbench (zimmer-tpami21a), rbv2_ranger (DBLP:conf/automl/PfistererSMBB22), PD1 (wang-jmlr24), and JAHS-Bench-201 (bansal-neurips22a). The implementation is based on shapiq (Muschalik.2024) and (will be) publicly available on GitHub11footnotemark: 1. We provide details regarding the setup, interpreting plots, and more results in the appendix. Generally, positive interactions are colored in red and negative in blue.
@@ 6.1 Insights from Ablation and Tunability
Figure 2: Left: Interaction graphs showing Möbius interactions (MI), second-order Shapley interactions (SI), and Shapley values (SV) where MIs terms are aggregated for interoperability. Right: Faithfulness of lower-order explanations approximating higher-order effects (Muschalik.2024). An explanation order of 3 already approximates the full game ($R^{2}\approx 1$) well.
Figure 3: Upset plots for Ablation (left) and Tunability (right) of lm1b_transformer (wang-jmlr24).
@@
First, we compare the results of the Ablation and the Tunability game in terms of hyperparameter importance and interactions (cf. Figure 3). We retrieve an optimized configuration of PD1’s lm1b_transformer scenario and explain it with the Ablation game. HyperSHAP’s explanation shows that the majority of the performance increase is attributed to the initial learning rate (L-I), which is not surprising since it is also intuitively the most important one. However, using HyperSHAP to create Tunability explanations reveals that both hyperparameters, L-I and optimizer momentum (O-M), are of equal importance with a negative interaction. Thus, the optimizer chose to tune L-I over O-M for the configuration in question, even though a similar performance improvement could have been achieved by tuning O-M instead. Hence, HyperSHAP can reveal which hyperparameters were subject to optimization via the Ablation game, while the Tunability game emphasizes the potential contributions of hyperparameters and their interactions.
@@ 6.2 Higher-Order Interactions in HPO
Second, we investigate the interaction structure of HPO problems for individual and across datasets. In Figure 2, left (MI), and further in the appendix, we observe the presence of many higher-order interactions, which are difficult to interpret. The SIs (order 2) and SV in HyperSHAP summarize the MI into interpretable explanations.
Figure 2, right, shows that SIs still faithfully capture the overall game behavior, which we measure with a Shapley-weighted loss (Muschalik.2024) and varying explanation order. We find that most of the explanatory power is captured by interactions up to the third order, confirming prior research that suggests hyperparameter interactions are typically of lower order (DBLP:conf/gecco/PushakH20a). Interactions beyond the third order contribute little to the overall understanding of the game. Thanks to the convenient properties of the SV and SIs, HyperSHAP provides a reliable way to capture and fairly summarize higher-order interactions into more interpretable explanations.
Figure 5: Explaining the surrogate model in SMAC’s Bayesian optimization with MIs at 1%, 5%, 25%, and 100% of the budget. Over time, SMAC notices first the importance of N-L and later B-S.
@@ 6.3 Detecting Optimizer Bias
The third experiment uses the Optimizer Bias game to investigate biases in black-box hyperparameter optimizers. To this end, we create two artificially biased hyperparameter optimizers. The first optimizer tunes each hyperparameter separately, ignoring interactions between them, while the second is not allowed to tune the most important hyperparameter. The virtual best hyperparameter optimizer is an ensemble of the investigated optimizer and five random search optimizers with a budget of 10,000 samples each. Ideally, a perfect optimizer would show no interactions and no main effects in HyperSHAP’s Optimizer Bias explanations as the differences for every coalition would be 0.
@@
Fig. 4 shows the Optimizer Bias explanations, i.e., the difference between two Tunability games, using the optimizer’s returned value and the (approximated) maximum, respectively. Note that main effects in the Optimizer Bias game can only be negative and show the optimizer’s inability to properly tune certain hyperparameters. In Fig. 4a, small main effects, in turn, suggest that the optimizer can effectively tune hyperparameters individually. The presence of both negative and positive interactions, which result from missing out on positive and negative interactions, respectively, shows that it fails to capture interactions. This confirms that the optimizer, which tunes hyperparameters independently, fails to capture their joint synergies. On the other hand, the second optimizer, ignoring the weight decay (W-D) hyperparameter for this particular dataset, clearly demonstrates bias in the interaction graph in Fig. 4b. The blue main effect for W-D and interactions involving W-D reveal this bias, showing how HyperSHAP can help identify such flaws and contribute to the development of more effective HPO methods.
6.4 Explaining Bayesian Optimization
Inspired by LMU, we use HyperSHAP to explain SMAC (lindauer-jmlr22a) – a state-of-the-art hyperparameter optimizer based on Bayesian optimization – by analyzing its surrogate model throughout the optimization process. We run SMAC with a total budget of 6 0006\,000 evaluations and inspect the surrogate model at $1\%,5\%,25\%,$ and $100\%$ of the budget. As shown in Fig. 5, we observe how the model’s belief about the main effects of and interactions between hyperparameters evolves over time.
Early in the process, the surrogate model exhibits numerous large higher-order interactions, reflecting high uncertainty and a broad range of plausible interactions. Despite this, it already identifies N-L as an important hyperparameter. As optimization progresses, the model’s uncertainty decreases, leading to lower interaction values. By 25% of the budget, the optimizer uncovers B-S as another important hyperparameter, and the surrogate model briefly broadens its hypothesis about the performance landscape, resulting in increased interaction values. Eventually, as SMAC converges, the model refines its understanding of the performance landscape and interactions that are considered plausible are reduced. This analysis illustrates how HyperSHAP can provide insights into the evolving dynamics of HPO processes.
6.5 Comparison with fANOVA
@@
Figure 6: Anytime performance plots of HPO runs involving only the top-2 important hyperparameters for two datasets of lcbench (zimmer-tpami21a).
Lastly, considering the experimental setting proposed with fANOVA by fANOVA, we compare HyperSHAP and fANOVA on the task of selecting a subset of important hyperparameters to tune. To this end, for a given HPO task, we run fANOVA and HyperSHAP using the SV with the Tunability and Sensitivity game to obtain explanations of order 1. Selecting the two most important hyperparameters, we conduct a subsequent HPO run optimizing only these two hyperparameters. Since Tunability quantifies performance gains (tunability), we expect explanations of this method to be more suitable for this specific task than explanations from the Sensitivity game, which quantifies variances in performances while tuning a certain hyperparameter. Also, the importance quantified over all variances, as in fANOVA, could yield less actionable explanations.
The results in Fig. 6 confirm that the anytime performance of the runs informed by HyperSHAP is superior to that informed by fANOVA, and that Tunability outperforms Sensitivity here. For this specific task, the explanations from the Tunability game are more suitable.
6.6 Runtime Analysis
We found HyperSHAP to be efficient across all evaluated settings. An Ablation game took 5s to 2m, while Tunability ranged from 6m to 15m for 7 and 4 hyperparameters, respectively, showing high dependency on the surrogate’s efficiency, and up to 8.5h for 10 hyperparameters on the JAHS-Bench-201 benchmark. As all coalition evaluations are independent, the method is highly parallelizable, enabling substantial wall-clock reductions. Overall, this adds modest overhead to typical multi-hour HPO runs.
@@ 7 Conclusion
In this paper, we proposed HyperSHAP, a post-hoc explanation framework for consistently and uniformly explaining hyperparameter optimization using the SV and SIs across three levels: hyperparameter values, sensitivity and tunability of learners, and optimizer capabilities. Unlike previous methods that quantify variance (fANOVA), HyperSHAP attributes performance contributions. We demonstrated that HyperSHAP not only enhances understanding of the impact of hyperparameter values or tunability of learners but also provides actionable insights for optimizing hyperparameters on related tasks.
The computational bottleneck is the approximation of the $\max$ over $\bm{\lambda}\in\bm{\Lambda}$ via simulated HPO, requiring research on more efficient yet unbiased methods, e.g., via Bayesian algorithm execution (bobax). Furthermore, extensions of HyperSHAP to the analysis of optimizing machine learning pipelines are important future work (HeffetzVKR20; FeurerEFLH22; wever-ida2020a; wever-tpami21a). Additionally, we plan to develop HPO methods that use HPI to learn across datasets for improving their efficiency. This may allow warm-starting HPO in an interpretable way, complementing recent work on prior-guided HPO (HvarfnerHN24; fehring-arxiv25) and human-centered AutoML (LindauerKKMT0HF24).
Acknowledgements
Marcel Wever and Marius Lindauer gratefully acknowledge funding by the European Union (ERC, “ixAutoML”, grant no.101041029). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Maximilian Muschalik gratefully acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation): TRR 318/1 2021 – 438445824.
Organization of the Appendix
The technical supplement is organized as follows.
@@ A.1 Proof of Proposition 1
$\nu(S)$ $=\max_{\bm{\lambda}\in\bm{\Lambda}}\,\textsc{Val}_{u}(\bm{\lambda}\oplus_{S}\bm{\lambda}^{0},D)$
$=\max_{\bm{\lambda}^{\ast}\in A}\textsc{Val}_{u}(\bm{\lambda}^{\ast},D)\overset{A\subseteq B}{\leq}\max_{\bm{\lambda}^{\ast}\in B}\textsc{Val}_{u}(\bm{\lambda}^{\ast},D)$
$=\max_{\bm{\lambda}\in\bm{\Lambda}}\,\textsc{Val}_{u}(\bm{\lambda}\oplus_{T}\bm{\lambda}^{0},D)$
$=\nu(T).$
This concludes that the Tunability game is monotone. As a consequence, we obtain non-negative SVs due to the monotonicity axiom (Fujimoto.2006) of the SV. We can also give a direct proof of this via the well-known representation of the SV in terms of a weighted average over marginal contributions as
@@
Moreover, the pure individual (main) effects obtained from the functional ANOVA framework are represented by the MI of the individuals (Fumagalli.2024a). By the monotonicity of $\nu$, we obtain again
@@ A.2 Proof for Table 1
Given the synthetic Tunability and Sensitivity game with two dimensions $\mathcal{N}=\{1,2\}$, our goal is to show that the main and interaction effects are given by Tab. 1.
@@
We first proceed to compute the game values of the Tunability game for $S\subseteq\mathcal{N}$ with the optimal configuration $\bm{\lambda}^{\ast}=(1,m)$ as
$\nu_{G_{T}}(S)$ $=\max_{\bm{\lambda}\in\bm{\Lambda}}\textsc{Val}_{u}(\bm{\lambda}\oplus_{S}\bm{\lambda}^{0},D)$ (1)
$=\textsc{Val}_{u}(\bm{\lambda}^{*}\oplus_{S}\bm{\lambda}^{0},D)$ (2)
$=\sum_{i\in\{1,2\}}\mathbf{1}_{(\bm{\lambda}\oplus_{S}\bm{\lambda}^{0})_{i}=\lambda^{\ast}_{i}}$ (3)
$=|S|+\sum_{i\in\mathcal{N}:i\notin S}\mathbf{1}_{\lambda^{0}_{i}=\lambda^{\ast}_{i}}.$ (4)
For the baseline set to $\bm{\lambda}^{0}:=(0,0)$, the second sum in Eq. 1 vanishes and we thus obtain
@@
Clearly, the interaction $\lambda_{1}\times\lambda_{2}$, i.e., $m(\{1,2\})$, vanishes, and thus the HPI scores of the individuals are given by their main effects in terms of the MIs. In summary, the HPI main effects using the SV and the MI are both equal to $1$, whereas the interaction is zero, confirming the values shown in Tab. 1.
For the baseline set to $\bm{\lambda}^{0}:=\bm{\lambda}^{\ast}$, the second sum in Eq. 1 equals $|\mathcal{N}|-|S|$ and thus we obtain a constant game
@@
Consequently, all interactions and main effects are zero due to the dummy axiom (Fujimoto.2006), confirming Tab. 1.
@@
$\nu_{G_{V}}(S)$ $=\mathbb{V}_{\bm{\lambda}^{\ast}\sim p(\bm{\lambda}^{\ast})}[\textsc{Val}_{u}(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0},D)]$
$=\mathbb{V}_{\bm{\lambda}^{\ast}\sim p(\bm{\lambda}^{\ast})}[\sum_{i\in\{1,2\}}\mathbf{1}_{(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0})_{i}=\lambda^{0}_{i}}]$
$=\sum_{i\in S}\mathbb{V}_{\lambda_{i}^{\ast}\sim p(\lambda_{i}^{\ast})}[\mathbf{1}_{(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0})_{i}=\lambda^{0}_{i}}]$
@@
Given any baseline, we have $\mathbf{1}_{(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0})_{i}=\lambda^{0}_{i}}\sim\text{Ber}(q_{i})$ with $q_{1}=1/2$ and $q_{2}=1/(m+1)$ due to the uniform distribution, which sets this value to $1$, if the optimal configuration value is chosen. The variance of this Bernoulli variable is then given by q(1q)q(1-q), which yields
@@
[…]
@@
which confirms the values given in Tab. 1 and concludes the proof. ∎
A.3 Proof of Theorem 1
Proof.
Consider an upper bound for the approximation $|\hat{\nu}(T)-\nu(T)|\leq\epsilon$ for all $T\subseteq N$. We can then naively bound the difference of Shapley values using its representation over marginal contributions as follows:
$|\phi_{\hat{\nu}}(i)-\phi_{\nu}(i)|\leq\sum_{T\subseteq N\setminus i}\frac{1}{n\binom{n-1}{t}}|\hat{\nu}(T\cup i)-\nu(T\cup i)-(\hat{\nu}(T)-\nu(T))|\leq\sum_{T\subseteq N\setminus i}\frac{1}{n\binom{n-1}{t}}2\epsilon=2\epsilon\,$,
where we have used for the sum that the Shapley value is a probabilistic value. A similar bound can be established for order-$k$ Shapley interactions by bounding their discrete derivatives with $2^{k}\epsilon$. ∎
A.4 Example: Non-Monotone Sensitivity Game
In this section, we give an example of a non-monotone Sensitivity game. To this end, we consider two hyperparameters $\mathcal{N}=\{1,2\}$ equipped with independent Bernoulli distributions $\lambda_{1},\lambda_{2}\overset{\text{iid}}{\sim}\text{Ber}(1/2)$. We consider a performance measure as